OpenStudy (anonymous):

can anyone help me write an equation of a transformed graph of tangent with period 4 that has been shifted vertically up 3 units

OpenStudy (jdoe0001):

Atan( Bx + C ) + D D = vertical shift, D > 0, up, D < 0, down B = parent function's original period change by \(\bf \cfrac{\text{original period}}{B}\) tangent's original period is \(\bf \large \pi\)

OpenStudy (anonymous):

so the equation would be....

OpenStudy (anonymous):

tan(4x)+3?

OpenStudy (jdoe0001):

well, you want the original period to change to 4 that means you want \(\bf \cfrac{\text{original period}}{B} = 4 \implies \cfrac{\pi}{B}=4\) so solve for "B" to see which B you can use to make the period to 4

OpenStudy (jdoe0001):

the +3 part is ok, yes, that'd shift it upwards by 3 units

OpenStudy (anonymous):

than is it 1/4?

OpenStudy (jdoe0001):

yes well, pi/4 that is \(\bf \cfrac{\pi}{B}=4 \implies B = \cfrac{\pi}{4}\\ tan\left(\cfrac{\pi}{4}x\right)+3\)